2720 people attended the first home football game of the season. Student tickets were $3 and non-student tickets were $5. The total revenue for tickets was $11,390. How many students attended the game?
Answer by checkley75(3666)

S+NS=2720 S=2720-NS
3S+5NS=11390
NOW SUBSTITUTE (2720-NS)FOR S IN THIS EQUATION & SOLVE FOR NS
3(2720-NS)+5NS=11390
8160-3NS+5NS=11390
2NS=11390-8160
2NS=3230
NS=3230/2
NS=1615 NUMBER OF NON STUDENT TICKETS SOLD.
S+1615=2720
S=2720-1615
S=1105 STUDENT TICKETS SOLD.
PROOF
3*1105+5*1615=11390
3315+8075=11390
11390=11390